OpenStudy (anonymous):

Solve x – 2y = 6 for y.

OpenStudy (jiteshmeghwal9):

(i) subtract x from both sides (ii)then divide -2 from both sides

OpenStudy (anonymous):

thanks

OpenStudy (jiteshmeghwal9):

yw :)

Parth (parthkohli):

Solve for a variable by taking all its friends away.

Parth (parthkohli):

Example! In \(\rm a + b = 9\), \(\rm a\) and \(\rm b\) are friends, and you have to solve for \(\rm a\). For that, you gotta take its friend \(\rm b\) away. Take \(\rm b\) to the other side. See that \(\rm b\) was happy(positive), but now it gets angry(negative). This becomes\[\rm a = 9 - b\]

OpenStudy (anonymous):

x - 2y = 6 for y -2y = 6 - x y = (6 - x)/-2?

Parth (parthkohli):

Yeah